OpenStudy (anonymous):

yay. pahelp nga po.. Find the inverse function of x= (4-3y)^(3/2)

OpenStudy (anonymous):

The function can be written like this : \[\Large f(y)=(4-3y)^{\frac32}=\sqrt{(4-3y)^3}\] To find the inverse function, we solve the equation f(y)=x. We have : \[\Large f(y)=x\iff \sqrt{(4-3y)^3}=x\iff (4-3y)^3=x^2\] \[\Large ~~\iff 4-3y=\sqrt[3]{x^2}\iff 3y=4-\sqrt[3]{x^2}\iff y=\frac{4-\sqrt[3]{x^2}}{3}\] So, the inverse function is : \[\Large f^{-1}(x)=\frac{4-\sqrt[3]{x^2}}{3}\]